OpenStudy (anonymous):

Help :c The navigator has to program the ship’s computer so that it will be prepared to intercept the satellite. To program it, the satelite’s path is set equal to the ship’s. The programmed equation is –4x – 2 = –x + 3. Explain to the navigator how the programmed equation can be graphed to find the point of intersection. Use complete sentences.

OpenStudy (anonymous):

@ganeshie8 help? :c

OpenStudy (anonymous):

@hartnn can you help me please?

OpenStudy (anonymous):

I know I have to put it into slope intercept form I keep getting it messed up though

OpenStudy (anonymous):

@mathslover can you help me please?

OpenStudy (anonymous):

@ganeshie8 I need your smartness ;-;

ganeshie8 (ganeshie8):

\(–4x – 2 = –x + 3\)

ganeshie8 (ganeshie8):

basically you want to solve it by graphing instead of algebraically

OpenStudy (anonymous):

I started by adding x to both sides

ganeshie8 (ganeshie8):

lol dont do that, the question is about graphing and solving. no algebra ok ?

OpenStudy (anonymous):

Okay .-.

ganeshie8 (ganeshie8):

\(–4x–2=–x+3 \) split it into two equations : \(y = -4x-2\) \(y = -x+3\) graph them in geogebra and find the intersection point.

OpenStudy (anonymous):

can I just use geogebra for that or do I solve more?

OpenStudy (anonymous):

oh wow I just read that, sorry

ganeshie8 (ganeshie8):

:)

OpenStudy (anonymous):

I have to redownload it, one sec

OpenStudy (anonymous):

okay it looks like for y=-4x-2 the point are -.5 and -2

ganeshie8 (ganeshie8):

try this online version : http://www.geogebra.org/webstart/geogebra.html

OpenStudy (anonymous):

I can't use the online version, this computer is weird, but I did the other one and the points were 3,3

ganeshie8 (ganeshie8):

hold up i plotted wrong equations

OpenStudy (anonymous):

I put the same thing but mine i different than yours. weird.

OpenStudy (anonymous):

okay

ganeshie8 (ganeshie8):

see if this looks right

ganeshie8 (ganeshie8):

the point of intersection is : \((-1.67, ~4.67)\)

OpenStudy (anonymous):

Yes, that's what I have, I was just confused on what you were asking for

OpenStudy (anonymous):

Thanks again! c:

ganeshie8 (ganeshie8):

np :)